OpenStudy (anonymous):

2^3 means we multiply 2 for 3 times but why is 2^0=1 whats the logic behind it and what is 0^0=?

OpenStudy (anonymous):

no it s not multiplying It's to say the power of a number\[x ^{y}\] like this

OpenStudy (anonymous):

i thing both are same 2^3=2*2*2=8 i.e 2 is multiplying 3 times

OpenStudy (anonymous):

2^3 = 2*2*2 so , this is called the two to power 3 but 2^0 always =1 any number to power 0 = 1

OpenStudy (anonymous):

0^0= 1 also or it is zero=???

OpenStudy (phi):

0^0 is a strange beast. Treat it as indeterminate, but see http://en.wikipedia.org/wiki/Exponentiation#Zero_to_the_zero_power, where there are cases where it is treated as 1. Also, I have attached a file about x^0